Circular No. 8860

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2007cr–2007cv Further to IAUC 8851, additional apparent supernovae have been dis-

covered on CCD frames: 2007cr and 2007cv by A. Klotz (0.25-m ‘TAROT’ robotic telescope, La Silla; R-band indicated), and the rest by W. Li (LOSS, unfiltered KAIT; with D. Winslow for 2007cs, and with N. Joubert for the rest). Discovery observations for objects brighter than mag 17:

SN  2007 UT  α_{2000}  δ_{2000}  Mag.  Offset
2007cr June 16.30 16^{h}21^{m}56.^{s}5 −60^{o}40′24′′ 15.0 20′′ E, 15′′ S
2007cs June 24.49 23 49 39.00 +29 55 52.1 15.9 12′′.4 W, 2′′.5 S
2007cv June 29.00 10 37 45.4 −27 04 57 15.3 24′′ W, 6′′ S

Additional information is available on Central Bureau Electronic Telegrams: 2007cr in ESO 137-19, CBET 985; 2007cs in UGC 12798, CBETs 986 and 987 (type-Ia, discovered some two weeks after maximum); 2007ct in NGC 6944A, CBETs 988 and 994 (mag 18.1 on June 28; type-II); 2007cu in UGC 10214, CBETs 988 and 994 (mag 18.9 on June 27; type-II); 2007cv in IC 2597, CBET 989.

COMETS 189P/NEAT AND 190P/MUELLER Comet P/2007 N2 = P/2002 O5 (cf. IAUC 8856) has been given the

permanent number 189P, and comet P/2007 O2 = P/1998 U2 (cf. IAUC 8859) has been given the permanent number 190P (cf. MPC 60237).

COMET P/2007 N1 (McNAUGHT) Additional observations have shown this comet (cf. IAUC 8855) to be

of short orbital period, as indicated by the following orbital elements (from MPC 60281):

T = 2007 Sept.15.892 TT  ω = 275.^{o}699
e = 0.38791  Ω = 105.867 }2000.0
q = 2.03021 AU  i = 8.721
a = 3.31687 AU n^{o} = 0.163159 P = 6.04 years

SATURN XLIV (HYRROKKIN) The IAU Working Group on Planetary Science Nomenclature has voted

to change the spelling of the name of Saturn XLIV from ‘Hyrokkin’ (as announced on IAUC 8826) to ‘Hyrrokkin’.

2007 July 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green